1. Differentiate the following functions:

(a) y = 4x3 + 5x-4 + sin 3x


1
Expert's answer
2021-05-07T10:10:19-0400

Solution.


"y=4x^3+5x-4+\sin 3x,\newline
y'=12x^2+5+3\cos 3x."